# Baichuan Alignment Technical Report

### v2

@@ 5.3.1 CFBench
We introduce CFBench[99], a large-scale Comprehensive Constraints Following Benchmark designed to evaluate the constraint adherence capabilities of LLMs. CFBench features 1,000 meticulously curated samples covering over 200 real-world scenarios and more than 50 NLP tasks. It systematically compiles constraints from authentic instructions and constructs an innovative framework for constraint types, comprising 10 primary categories and over 25 subcategories, ensuring seamless integration of each constraint within the instructions. The 10 primary categories, partially illustrated in Figure 6, are: Content, Numerical, Stylistic, Format, Linguistic, Situation, Example, Inverse, Contradictory, and Rule Constraints. To ensure that the evaluation of LLM outputs aligns with user perceptions, we propose an advanced methodology that integrates multi-dimensional assessment criteria with requirement prioritization. This approach covers various aspects of constraints, instructions, and requirement fulfillment, corresponding to the CSR, ISR, and PSR metrics, respectively.
@@
Table 4 presents the comprehensive evaluation results of CFBench and its subsets for the leading models, assessed using three key metrics. Overall, GPT-4o consistently outperforms across all splits and metrics, with Claude-3.5-Sonnet following closely. Even among the top-performing models, there are varying degrees of differentiation, highlighting CFBench’s robust capability to distinguish the constraint-following proficiency of LLMs. Notably, Baichuan-Instruct demonstrates exceptional overall performance, exhibiting strong competitiveness among its peers, particularly excelling on the Hard Set. This success is largely attributed to the constraint expansion methods detailed in Section 4.1, Constraint Instruction Construction.
@@ 5.3.2 SysBench
Table 5 shows the CSR metric evaluation results in SysBench for several leading LLMs. GPT-4o and Claude 3.5 rank first and second in the Total Score, respectively. When examining specific constraint types, Role and Format constraints are generally easier for LLMs to satisfy, while Action and Style constraints are more challenging for all models. Besides GPT-4o and Claude 3.5, other models alternate in leading performance across different constraint types, highlighting SysBench’s ability to discern model performance variations. Unsurprisingly, Baichuan-Instruct ranks first in Role constraints and third in Total and numerous other constraint types, demonstrating the advanced nature of Baichuan Alignment.
@@ 5.3.3 FB-Bench
The subset evaluation results of leading LLMs in FB-Bench are presented in Table 6. Most LLMs exhibit superior performance in error correction compared to response maintenance. This suggests that most LLMs lack a robust capacity to differentiate between valid and misleading instructions. This deficiency could stem from the fact that optimizing an LLM’s instruction-following ability is relatively straightforward; however, overly adherent instruction-following can cause the LLM to distort reality. For error correction, Claude-3.5-Sonnet significantly outperform other LLMs. Conversely, in response to maintenance scenarios, ERNIE-4-8K, Qwen2-72B-Instruct and Baichuan-Instruct demonstrate competitiveness among peers.
@@ 6 Conclusion
In this report, we provide a comprehensive overview of Baichuan Alignment, an advanced, precise, and versatile alignment technique that has been widely applied across the Baichuan series of models. Our discussion covers various aspects, including optimization, data, key capability enhancements, and comprehensive evaluations. We cover the three stages of prompt augmentation system (PAS), supervised fine-tuning (SFT), and preference alignment, documenting and summarizing the challenges encountered, successful experiences, and some erroneous attempts made by the Baichuan Alignment team during the alignment phase. Through multi-dimensional evaluations of models optimized via Baichuan Alignment, including user experience, open-source benchmarks, and specific capability assessments, we demonstrate the effectiveness of Baichuan Alignment in addressing diverse scenario-specific issues and its adaptability to various base models. This report represents the industry’s inaugural comprehensive exposition of alignment technology to the public. We aspire for our methodologies, insights, and experiences to resonate with and inspire the research community, providing valuable guidance and lessons from our successes and missteps. Our ultimate goal is to contribute to the advancement of artificial intelligence and to collectively advance towards the realization of Artificial General Intelligence (AGI).

### v3

@@ 5.3.1 CFBench
We introduce CFBench[99], a large-scale Comprehensive Constraints Following Benchmark designed to evaluate the constraint adherence capabilities of LLMs. CFBench features 1,000 meticulously curated samples covering over 200 real-world scenarios and more than 50 NLP tasks. It systematically compiles constraints from authentic instructions and constructs an innovative framework for constraint types, comprising 10 primary categories and over 25 subcategories, ensuring seamless integration of each constraint within the instructions. The 10 primary categories include Content, Numerical, Stylistic, Format, Linguistic, Situation, Example, Inverse, Contradictory, and Rule Constraints. To ensure that the evaluation of LLM outputs aligns with user perceptions, we propose an advanced methodology that integrates multi-dimensional assessment criteria with requirement prioritization. This approach covers various aspects of constraints, instructions, and requirement fulfillment, corresponding to the CSR, ISR, and PSR metrics, respectively.
@@
Table 4 presents the comprehensive evaluation results of CFBench and its subsets for the leading models, assessed using three key metrics. Overall, GPT-4o consistently outperforms across all splits and metrics, with Claude-3.5-Sonnet following closely. Even among the top-performing models, there are varying degrees of differentiation, highlighting CFBench’s robust capability to distinguish the constraint-following proficiency of LLMs. Notably, Qwen2-Nova-72B demonstrates exceptional overall performance, exhibiting strong competitiveness among its peers, particularly excelling on the Hard Set. This success is largely attributed to the constraint expansion methods detailed in Section 4.1, Constraint Instruction Construction.
@@ 5.3.2 SysBench
Table 5 shows the CSR metric evaluation results in SysBench for several leading LLMs. GPT-4o and Claude 3.5 rank first and second in the Total Score, respectively. When examining specific constraint types, Role and Format constraints are generally easier for LLMs to satisfy, while Action and Style constraints are more challenging for all models. Besides GPT-4o and Claude 3.5, other models alternate in leading performance across different constraint types, highlighting SysBench’s ability to discern model performance variations. Unsurprisingly, Qwen2-Nova-72B ranks first in Role constraints, third in Total constraints, and performs highly in various other constraint types, demonstrating the advanced nature of Nova Alignment.
@@ 5.3.3 FB-Bench
The subset evaluation results of leading LLMs in FB-Bench are presented in Table 6. Most LLMs exhibit superior performance in error correction compared to response maintenance. This suggests that most LLMs lack a robust capacity to differentiate between valid and misleading instructions. This deficiency could stem from the fact that optimizing an LLM’s instruction-following ability is relatively straightforward; however, overly adherent instruction-following can cause the LLM to distort reality. For error correction, Claude-3.5-Sonnet significantly outperform other LLMs. Conversely, in response to maintenance scenarios, ERNIE-4-8K, Qwen2-72B-Instruct and Qwen2-Nova-72B demonstrate competitiveness among peers.
@@ 6 Conclusion
In this report, we provide a comprehensive overview of Nova Alignment, an advanced, precise, and versatile alignment technique that has been widely applied and validated across a range of outstanding models. Our discussion covers various aspects, including optimization, data, key capability enhancements, and comprehensive evaluations. We cover the three stages of prompt augmentation system (PAS), supervised fine-tuning (SFT), and preference alignment, documenting and summarizing the challenges encountered, successful experiences, and some erroneous attempts made by the Nova Alignment team during the alignment phase. Through multi-dimensional evaluations of models optimized via Nova Alignment, including user experience, open-source benchmarks, and specific capability assessments, we demonstrate the effectiveness of Nova Alignment in addressing diverse scenario-specific issues and its adaptability to various base models. This report represents the industry’s inaugural comprehensive exposition of alignment technology to the public. We aspire for our methodologies, insights, and experiences to resonate with and inspire the research community, providing valuable guidance and lessons from our successes and missteps. Our ultimate goal is to contribute to the advancement of artificial intelligence and to collectively advance towards the realization of Artificial General Intelligence (AGI).